OpenStudy (anonymous):

in a cricket match the batsman is ata point O and he hits the ball with velocity (7i + 24j) m/s. a fielder at point A position vector 2i + 3j relative to O, does not move at all as the ball passes by. assuming the ball suffers no change in velocity:

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

find the least distance between the fielder at point A and the ball during is journey to the boundary...

OpenStudy (anonymous):

umm sorry i dont have any idea because we dont know the time the velocity is \[\sqrt{7^{2}+24^{2}}=25\] i dont have any idea whats the next step

OpenStudy (anonymous):

part a of the wuestion said how long will the ball take to reach the boundary, 60 m from O. i got 2.4 seconds? help anymore?

OpenStudy (anonymous):

question

OpenStudy (anonymous):

if the boundary is 60m from 0 then t = 60/25 = 2.4

OpenStudy (nikvist):

\[l^2=(7t-2)^2+(24t-3)^2\] \[2l\frac{dl}{dt}=14(7t-2)+48(24t-3)=0\] \[98t-28+1152t-144=0\] \[1250t=172\] \[t=0.1376s\] \[l_{min}=\sqrt{(7\cdot 0.1376-2)^2+(24\cdot 0.1376-3)^2}\] \[l_{min}=1.08m\]

OpenStudy (anonymous):

how did you get l^2=(7t−2)^2+(24t−3)^2 ?

OpenStudy (nikvist):

least distance between two point (7t,24t) and (2,3)